Holding period and annual​ (investment) returns. Bohenick Classic Automobiles restores and rebuilds old classic cars. The company purchased and restored a classic 1957 Thunderbird convertible 5 years ago for $7,400.00. Today at​ auction, the car sold for $59,200.00. What are the holding period return and the annual return on this​ investment?

1. What is the holding period return of the​ car?

2. What is the simple annual return?

3. What is the EAR?
Answer #1

1

HPR = ending price/beginning price-1 = 59200/7400-1 = 700%

2

Simple return = HPR/years = 700/5 = 140%

3

EAR = (1+HPR)^(1/n)-1 = (1+7)^(1/5)-1 = 51.57%

=